# Situationally-Aware Dynamics Learning

### v1

@@ I Introduction
For reliable robot deployment in unstructured and dynamic environments, it is crucial for robotic systems to adapt to unforeseen challenges and operate effectively under uncertainty. Consider, for example, an Unmanned Ground Vehicle (UGV) navigating steep and rugged terrain. Such a robot may become immobilized due to unexpected factors, ranging from unmodeled terrain properties to variations in its own internal dynamics. These factors are inherently difficult, if not impossible, to anticipate exhaustively [23, 53, 32, 84, 99, 97]. This challenge is further exacerbated by the impracticality of equipping robots with every conceivable sensor and analysis algorithm needed to fully capture the world’s and robot’s full state in all scenarios. Consequently, robots must learn to operate effectively using only partial information derived from their observable state—the limited set of measurements and estimates available in real time. This necessitates the development of adaptive mechanisms capable of inferring and representing the hidden aspects of the robot and its environment to inform decision-making and enhance resilience in uncertain conditions [65, 26, 48, 45].
@@
Furthermore, when a robot relies on partial information to operate on a complex environment, the risk of ambiguous scenario representation becomes significant [66, 73, 12]. Such ambiguity is undesirable, as it increases uncertainty during decision-making, often leading to suboptimal plans and, in some cases, unsafe or hazardous behaviors. This challenge is exemplified in Figure 1: a UGV with incomplete state information—lacking a (robust) scene interpretation module to detect the terrain type it is traversing—executes an action from what appears to be the same state. However, the outcomes of these actions differ drastically, revealing a discrepancy that cannot be attributed solely to aleatoric errors (those arising from process or observation noise). This highlights the critical need for robust methods to infer and resolve hidden state ambiguities to ensure reliable robot operation.
@@
To address this, we propose a framework for learning online a representation of the unmodeled and unobserved latent factors that cause such ambiguities. This represents a significant challenge, as it requires balancing the tradeoff between representation capacity, data- and learning-efficiency, and minimization of inductive biases that could inadvertently exclude critical latent factors [57, 39]. Considering this, we introduce an online hidden state representation learning method that enables the robot to recover and leverage representations of latent world and robot factors in real time. By utilizing this learned representation, the robot can dynamically adapt its behavior and reliably accomplish its tasks, even in uncertain and unstructured environments.
@@
Recent studies have introduced various methods that leverage learned auxiliary representations during dynamics or policy learning to enable robot adaptation [58, 50, 51, 74, 61]. However, these methods are data-intensive and often rely on multi-stage procedures that require privileged information, such as ground-truth data that is difficult to obtain or is only available in simulation. To overcome these limitations, our proposed method learns such representations in an online and unsupervised manner, thus eliminating the need for multi-stage approaches. We achieved this by formulating the problem within the framework of Generalized Hidden Parameter Markov Decision Processes (GHP-MDPs) [70, 22].
@@
An important component of our framework is to represent the hidden state by learning online the joint distribution of the state transitions. We refer to this state transition distribution at a given moment as the situation of the robot. This approach enables the robot to capture an expressive probabilistic representation of its current context, encompassing both its state and the underlying environmental conditions. By leveraging proprioceptive sensing, such as inertial measurements, to assess the robot’s state, this representation becomes sufficiently general to effectively disambiguate the robot’s state across various scenarios. Additionally, as illustrated in Figure 1, the state evolution exhibits diverse probability density profiles under different conditions. This diversity facilitates the identification of latent factors that govern distinct scenarios. In this way, the situation can be understood as the Underlying Data Generating Process (UDGP)—a conceptual model that characterizes the true, often hidden, factors governing the system’s dynamics at any given moment [30, p. 109]. The UDGP encapsulates the combined influence of both robot-specific and environmental variables that affect the robot’s behavior, enabling a more accurate understanding of its state. To obtain this situation distribution in real-time as the robot operates, we introduce a multivariate extension of Bayesian Online Changepoint Detection (BOCD) [3]. This extension allows us to model the situation the robot is currently operating under and detect when a situation change occurs.
@@ II-A Neural-Symbolic Learning in Robotics
The field of neural-symbolic learning consists of two primary branches [93]. We will first focus on methods rooted in discrete logical reasoning and guiding abstractions. The second branch involves Physics-Infused Neural Networks (PINNs), which handle continuous abstractions to represent physical characteristics, which we cover later on.
Analogous to our work, Gopalan et al. [31] used changepoint detection for robotic applications. They use it to segment demonstration trajectories from which the robot learns skills. Then they learn symbols by identifying the termination conditions of the skills via clustering. Their objective was to map behaviors to natural language commands. However, for training they rely on demonstration trajectories that have been paired to natural language instructions, and find the symbols from the provided dataset. Conversely, in our case the agent discovers the symbols online and unsupervisedly, from the data generated by its proprioception, while it is exploring the environment.
[47] construct skill trees from demonstration trajectories. Each skill has a goal and an abstraction and they can later be composed. Similarly, they use a changepoint detection algorithm in charge of determining when the most relevant abstraction changes. However, they cannot recognize repeated skills and require the availability of an abstraction library while we directly model the robot’s dynamics which can then be used with sampling-based planners. Other authors have also worked along this line [49, 14, 68]. Moreover, learned symbolic abstractions for planning, skill learning and control have been explored in [82, 2, 8, 52, 76].
Research that is in proximity also includes Ly and Lipson [63] which propose a method for learning a system’s hybrid dynamics directly from time-series data through multi-modal symbolic regression. Similarly, [43] focus on estimating key physical parameters, such as tire slippage and longitudinal force, in the context of an autonomous vehicle model, offering critical insights for improving vehicle performance and control. Closer to our work, [104] employ physics-informed neural networks (PINNs) to predict motion in off-road driving scenarios, addressing the challenges posed by rough and unpredictable terrain while ensuring more accurate trajectory forecasting.
@@ II-B Model-Based Reinforcement Learning
System Identification (SI) aims to discover and model the relationship between the inputs and outputs of a system [5, 62]. MBRL is a SI method based on balancing exploration and exploitation. It iteratively collects more informative data for the task at hand, thanks to planning with an improved model learned with the data collected in previous iterations. Thus, the model evolves, leading to better planning performance which results in exploring more informative sections of the state-action space [89].
[1] explore how to use an approximate model with few real-life trials. They evaluate a policy using a sample trial from the real world, and leverage the approximate model to estimate the derivative of the evaluation with respect to the policy parameters. [21] introduce a model-based policy search strategy for high-dimensional control tasks. They employ Gaussian Processes to learn the dynamics and incorporate the model’s uncertainty during planning and policy evaluation. They highlight the data efficiency of this approach. Chua et al. [15] propose PETS, which uses ensembles of uncertainty-aware neural-networks together with sampling-based optimizers for planning.
Hansen et al. [37] presented TD-MPC2, which learns five components consisting of an (i) observation encoder, (ii) latent-space dynamics model, (iii) transition reward predictor, (iv) return or terminal value predictor, and (v) a policy predicting return-maximizing actions; to predict returns from a sequence of actions, usually from high-dimensional observations. Other works have also learned latent dynamics models [33, 35]. MoDem-V2 [55] builds on top of TD-MPC to learn vision-based models in the real-world for robot manipulation tasks. They identify unsafe exploration and over-optimism as a significant failure point for applying MBRL in the real-world and develop three techniques for safer real-world learning: policy centering, agency transfer and actor-critic ensembles.
@@ II-C1 Model-Free Reinforcement Learning:
[58] use a temporal convolutional network that takes a window of past proprioceptive states to generate control commands for quadruped locomotion over challenging terrains. This model is trained to “implicitly reason about contact and slippage events” from the proprioception. However, learning the previous model requires a teacher policy that has been previously obtained using privileged information of the terrain and the robot’s contact with it. Furthermore, they also make use of an adaptive curriculum in charge of modulating the terrain’s difficulty according the current performance of the policy. [50] proposed Rapid Motor Adaptation (RMA), which is a two-stage process to train adaptive policies via Model-Free RL (MFRL). First, an “environmental factors encoder” model is trained to map “privileged information” only available in simulation to a latent vector. This latent vector is used as input together with the current state and action to a “base-policy” trained via Proximal Policy Optimization [78, PPO]. Then, during the second phase, an “adaptation module” is trained to map a fixed-size window of past state-action pairs into a latent vector optimized to match the latent representations produced by the “environmental factors encoder”. This way, during deployment, the latent representation can be obtained without the need for privileged information. [51] add a third phase to the RMA framework to enable adaptive bipedal locomotion. The new stage involves using PPO to fine-tune the base policy but maintaining the adaptation module weights frozen, to ensure that the policy supports imperfect estimation of the latent representation. Follow up works to RMA include [74] which also uses proprioception history and privileged information to train a model for in-hand object rotation. Also, [61] learn the latent representation from depth perception for manipulation tasks.
@@ II-C2 Model-Based Reinforcement Learning:
[38] recognize that model-based methods are often restricted to tasks with limited dimensions or rely on certain assumptions about the system’s dynamics. Furthermore, they argue that high-dimensional models are challenging for planning due to increased compounding error, as also argued in [86, 90, 91]. Thus, they compress the state space to avoid learning unimportant features that could only contribute to compounding error during planning. Specifically, they propose to learn a state space encoder, a dynamics model in the latent space and action space, as well as a reward predictor. They place special emphasis in optimizing for reward prediction performance to encourage better compression of the state space. Closer to our work, [79] combine MBRL with primitives learned via MFRL to make model-based planning easier. In particular, they simultaneously identify “predictable behaviors” and the corresponding dynamics. Their key idea is to learn skills that are optimized to be easy to identify and for which a simpler accompanying dynamics model can be obtained. Finally, during deployment, the skill and corresponding dynamics model become a set of primitives that can be composed to solve the commanded task via model-based planning. In [80], it is shown that this framework can be leveraged for RL in the real world. [59] introduces Context-aware Dynamics Models (CaDM) to generalize across different dynamics. They learn a context encoder, forward and backward dynamics models. Both dynamics models conditioned by the latent vector in charge of encoding dynamics-specific information.
@@ II-C3 Situational Awareness:
Other methods that do not rely on RL but leverage representations of the system’s situation include pervasive computing, where situation identification is used to abstract noisy sensor data into high-level concepts [98]. For them, each situation is defined according to how the data is semantically interpreted by the users, which could require preliminary domain knowledge to obtain them. On the other hand, our work automatically builds situations by modelling the underlying data generation process that best explains the present sensor measurements. Langari and Won [56] use situation identification to determine a policy for intelligent energy management in parallel hybrid vehicles. They do it by abstracting environment, driving style and car operating mode information using short term statistical features. However, they learn to classify the situations in a supervised manner and selecting which statistics to use requires task-specific expert-knowledge, which we avoid by modeling the UDGP of the transition dynamics. Other robot applications involve human-robot-cooperation to determine whether the robot maintains SA based on surprise [19].
@@ II-D Generalized Hidden Parameter Markov Decision Process (GHP-MDP)
[70] introduce GHP-MDPs and develop a learning-based method where agents can generalize, in simulation, to the same family of tasks using transition models whose input is augmented with estimated latent factors. They model the latent factors with a diagonal Gaussian while minimizing the Kullback–Leibler divergence between the variational and prior distributions. The main difference to our approach, is that we detect the latent representation online, which allows us to directly augment the (state, action, next-state)-tuples for fine-tuning the dynamics with the symbol assigned to each representation of the transition distribution. Furthermore, we use a full-rank covariance matrix, instead of a less expressive diagonal covariance.
Related work also includes Hidden Parameter MDPs (HiP-MDPs) which assume the agent knows what it does not know. Specifically, it does not know the value of some critical environment variables and focuses on estimating these [22]. [103] learn state abstractions to enable multi-task RL. However, they assume the latent parameter is fixed for the duration of the episode, while in our case we are interested in detecting and accommodating to changes in the latent parameters during the episode. [42] extend this framework for robust and efficient transfer learning to subtle variations among tasks using Bayesian Neural Networks. [17] develop a method to express the transition probabilities in closed-form as a polynomial of the latent parameters, as well as maintaining a closed-form belief distribution of the latent parameters. They present promising results in two simulation tasks. In our case, the transition model’s input is augmented with the symbol of the detected latent representation, since we do not assume prior-knowledge about the variables that the agent will require to obtain a better model of the dynamics at deployment time.

### v2

@@ I Introduction
For reliable robot deployment in unstructured and dynamic environments, it is crucial for robotic systems to adapt to unforeseen challenges and operate effectively under uncertainty. Consider, for example, an Unmanned Ground Vehicle (UGV) navigating steep and rugged terrain. Such a robot may become immobilized due to unexpected factors, ranging from unmodeled terrain properties to variations in its own internal dynamics. These factors are inherently difficult, if not impossible, to anticipate exhaustively [du2011robot, kurniawati2012global, green2020planning, siva2021enhancing, yin2023reliable, xu2024boundary]. This challenge is further exacerbated by the impracticality of equipping robots with every conceivable sensor and analysis algorithm needed to fully capture the world’s and robot’s full state in all scenarios. Consequently, robots must learn to operate effectively using only partial information derived from their observable state—the limited set of measurements and estimates available in real time. This necessitates the development of adaptive mechanisms capable of inferring and representing the hidden aspects of the robot and its environment to inform decision-making and enhance resilience in uncertain conditions [mccallum1996hidden, ferguson2004pao, kroemer2014learning, kobayashi2022latent].
@@
Furthermore, when a robot relies on partial information to operate on a complex environment, the risk of ambiguous scenario representation becomes significant [meyer2011using, pramanick2022talk, celemin2023knowledge]. Such ambiguity is undesirable, as it increases uncertainty during decision-making, often leading to suboptimal plans and, in some cases, unsafe or hazardous behaviors. This challenge is exemplified in Figure 1: a UGV with incomplete state information—lacking a (robust) scene interpretation module to detect the terrain type it is traversing—executes an action from what appears to be the same state. However, the outcomes of these actions differ drastically, revealing a discrepancy that cannot be attributed solely to aleatoric errors (those arising from process or observation noise). This highlights the critical need for robust methods to infer and resolve hidden state ambiguities to ensure reliable robot operation.
@@
To address this, we propose a framework for learning online a representation of the unmodeled and unobserved latent factors that cause such ambiguities. This task is challenging because it requires balancing representation capacity, data- and learning-efficiency, and minimizing inductive biases that might inadvertently exclude critical latent factors [langdon2019uncovering, hundt2006representing]. Motivated by these constraints, we introduce an online hidden-state representation learning method that enables the robot to recover and leverage compact representations of latent world and robot factors in real time. By exploiting this learned representation the robot can dynamically adapt its behavior and reliably accomplish tasks even in uncertain and unstructured environments. Prior work has attempted to address this by learning distributions over causal structures—an approach explicitly intended to account for uncertainty in the dynamics model caused in part by unobserved latent factors [murillogonzalez2025causaldistributions]; our method instead targets compact, real-time latent state representations for online adaptation.
@@
Recent studies have introduced various methods that leverage learned auxiliary representations during dynamics or policy learning to enable robot adaptation [lee2020learning, kumar2021rma, kumar2022adapting, qi2023hand, liang2024rapid]. However, these methods are data-intensive and often rely on multi-stage procedures that require privileged information, such as ground-truth data that is difficult to obtain or is only available in simulation. To overcome these limitations, our proposed method learns such representations in an online and unsupervised manner, thus eliminating the need for multi-stage approaches. We achieved this by formulating the problem within the framework of Generalized Hidden Parameter Markov Decision Processes (GHP-MDPs) [perez2020generalized, doshi2016hidden].
@@
An important component of our framework is to represent the hidden state by learning online the joint distribution of the state transitions. We refer to this state transition distribution at a given moment as the situation of the robot. This approach enables the robot to capture an expressive probabilistic representation of its current context, encompassing both its state and the underlying environmental conditions. By leveraging proprioceptive sensing, such as inertial measurements, to assess the robot’s state, this representation becomes sufficiently general to effectively disambiguate the robot’s state across various scenarios. Additionally, as illustrated in Figure 1, the state evolution exhibits diverse probability density profiles under different conditions. This diversity facilitates the identification of latent factors that govern distinct scenarios. In this way, the situation can be understood as the Underlying Data Generating Process (UDGP)—a conceptual model that characterizes the true, often hidden, factors governing the system’s dynamics at any given moment [Goodfellow-et-al-2016, p. 109]. The UDGP encapsulates the combined influence of both robot-specific and environmental variables that affect the robot’s behavior, enabling a more accurate understanding of its state. To obtain this situation distribution in real-time as the robot operates, we introduce a multivariate extension of Bayesian Online Changepoint Detection (BOCD) [adams2007bayesian]. This extension allows us to model the situation the robot is currently operating under and detect when a situation change occurs.
@@ II-A Neural-Symbolic Learning in Robotics
The field of neural-symbolic learning consists of two primary branches [wang2024imperative]. We will first focus on methods rooted in discrete logical reasoning and guiding abstractions. The second branch involves Physics-Infused Neural Networks (PINNs), which handle continuous abstractions to represent physical characteristics, which we cover later on.
Analogous to our work, gopalan2020simultaneously used changepoint detection for robotic applications. They use it to segment demonstration trajectories from which the robot learns skills. Then they learn symbols by identifying the termination conditions of the skills via clustering. Their objective was to map behaviors to natural language commands. However, for training they rely on demonstration trajectories that have been paired to natural language instructions, and find the symbols from the provided dataset. Conversely, in our case the agent discovers the symbols online and unsupervisedly, from the data generated by its proprioception, while it is exploring the environment.
[konidaris2012robot] construct skill trees from demonstration trajectories. Each skill has a goal and an abstraction and they can later be composed. Similarly, they use a changepoint detection algorithm in charge of determining when the most relevant abstraction changes. However, they cannot recognize repeated skills and require the availability of an abstraction library while we directly model the robot’s dynamics which can then be used with sampling-based planners. Other authors have also worked along this line [kulic2009online, chiappa2010movement, niekum2012learning]. Moreover, learned symbolic abstractions for planning, skill learning and control have been explored in [silver2022learning, achterhold2023learning, belta2007symbolic, kumar2023learning, rosen2023synthesizing].
Research that is in proximity also includes ly2012learning which propose a method for learning a system’s hybrid dynamics directly from time-series data through multi-modal symbolic regression. Similarly, [kim2022physics] focus on estimating key physical parameters, such as tire slippage and longitudinal force, in the context of an autonomous vehicle model, offering critical insights for improving vehicle performance and control. Closer to our work, [zhao2024physord] employ physics-informed neural networks (PINNs) to predict motion in off-road driving scenarios, addressing the challenges posed by rough and unpredictable terrain while ensuring more accurate trajectory forecasting.
@@ II-B Model-Based Reinforcement Learning
System Identification (SI) aims to discover and model the relationship between the inputs and outputs of a system [aastrom1971system, ljung2010perspectives]. MBRL is a SI method based on balancing exploration and exploitation. It iteratively collects more informative data for the task at hand, thanks to planning with an improved model learned with the data collected in previous iterations. Thus, the model evolves, leading to better planning performance which results in exploring more informative sections of the state-action space [sutton2018reinforcement].
[abbeel2006using] explore how to use an approximate model with few real-life trials. They evaluate a policy using a sample trial from the real world, and leverage the approximate model to estimate the derivative of the evaluation with respect to the policy parameters. [deisenroth2011pilco] introduce a model-based policy search strategy for high-dimensional control tasks. They employ Gaussian Processes to learn the dynamics and incorporate the model’s uncertainty during planning and policy evaluation. They highlight the data efficiency of this approach. chua2018deep propose PETS, which uses ensembles of uncertainty-aware neural-networks together with sampling-based optimizers for planning.
hansen2023td presented TD-MPC2, which learns five components consisting of an (i) observation encoder, (ii) latent-space dynamics model, (iii) transition reward predictor, (iv) return or terminal value predictor, and (v) a policy predicting return-maximizing actions; to predict returns from a sequence of actions, usually from high-dimensional observations. Other works have also learned latent dynamics models [ha2018world, hafner2019dream]. MoDem-V2 [lancaster2023modem] builds on top of TD-MPC to learn vision-based models in the real-world for robot manipulation tasks. They identify unsafe exploration and over-optimism as a significant failure point for applying MBRL in the real-world and develop three techniques for safer real-world learning: policy centering, agency transfer and actor-critic ensembles.
@@ II-C1 Model-Free Reinforcement Learning:
[lee2020learning] use a temporal convolutional network that takes a window of past proprioceptive states to generate control commands for quadruped locomotion over challenging terrains. This model is trained to “implicitly reason about contact and slippage events” from the proprioception. However, learning the previous model requires a teacher policy that has been previously obtained using privileged information of the terrain and the robot’s contact with it. Furthermore, they also make use of an adaptive curriculum in charge of modulating the terrain’s difficulty according the current performance of the policy. [kumar2021rma] proposed Rapid Motor Adaptation (RMA), which is a two-stage process to train adaptive policies via Model-Free RL (MFRL). First, an “environmental factors encoder” model is trained to map “privileged information” only available in simulation to a latent vector. This latent vector is used as input together with the current state and action to a “base-policy” trained via Proximal Policy Optimization [schulman2017proximal, PPO]. Then, during the second phase, an “adaptation module” is trained to map a fixed-size window of past state-action pairs into a latent vector optimized to match the latent representations produced by the “environmental factors encoder”. This way, during deployment, the latent representation can be obtained without the need for privileged information. [kumar2022adapting] add a third phase to the RMA framework to enable adaptive bipedal locomotion. The new stage involves using PPO to fine-tune the base policy but maintaining the adaptation module weights frozen, to ensure that the policy supports imperfect estimation of the latent representation. Follow up works to RMA include [qi2023hand] which also uses proprioception history and privileged information to train a model for in-hand object rotation. Also, [liang2024rapid] learn the latent representation from depth perception for manipulation tasks.
@@ II-C2 Model-Based Reinforcement Learning:
[havens2019learning] recognize that model-based methods are often restricted to tasks with limited dimensions or rely on certain assumptions about the system’s dynamics. Furthermore, they argue that high-dimensional models are challenging for planning due to increased compounding error, as also argued in [stachowicz2024racer, talvitie2014model, talvitie2017self]. Thus, they compress the state space to avoid learning unimportant features that could only contribute to compounding error during planning. Specifically, they propose to learn a state space encoder, a dynamics model in the latent space and action space, as well as a reward predictor. They place special emphasis in optimizing for reward prediction performance to encourage better compression of the state space. Closer to our work, [sharma2019dynamics] combine MBRL with primitives learned via MFRL to make model-based planning easier. In particular, they simultaneously identify “predictable behaviors” and the corresponding dynamics. Their key idea is to learn skills that are optimized to be easy to identify and for which a simpler accompanying dynamics model can be obtained. Finally, during deployment, the skill and corresponding dynamics model become a set of primitives that can be composed to solve the commanded task via model-based planning. In [sharma2020emergent], it is shown that this framework can be leveraged for RL in the real world. [lee2020context] introduces Context-aware Dynamics Models (CaDM) to generalize across different dynamics. They learn a context encoder, forward and backward dynamics models. Both dynamics models conditioned by the latent vector in charge of encoding dynamics-specific information.
@@ II-C3 Situational Awareness:
Other methods that do not rely on RL but leverage representations of the system’s situation include pervasive computing, where situation identification is used to abstract noisy sensor data into high-level concepts [ye2012situation]. For them, each situation is defined according to how the data is semantically interpreted by the users, which could require preliminary domain knowledge to obtain them. On the other hand, our work automatically builds situations by modelling the underlying data generation process that best explains the present sensor measurements. langari2005intelligent use situation identification to determine a policy for intelligent energy management in parallel hybrid vehicles. They do it by abstracting environment, driving style and car operating mode information using short term statistical features. However, they learn to classify the situations in a supervised manner and selecting which statistics to use requires task-specific expert-knowledge, which we avoid by modeling the UDGP of the transition dynamics. Other robot applications involve human-robot-cooperation to determine whether the robot maintains SA based on surprise [dahn2018situation].
@@ II-D Generalized Hidden Parameter Markov Decision Process (GHP-MDP)
[perez2020generalized] introduce GHP-MDPs and develop a learning-based method where agents can generalize, in simulation, to the same family of tasks using transition models whose input is augmented with estimated latent factors. They model the latent factors with a diagonal Gaussian while minimizing the Kullback–Leibler divergence between the variational and prior distributions. The main difference to our approach, is that we detect the latent representation online, which allows us to directly augment the (state, action, next-state)-tuples for fine-tuning the dynamics with the symbol assigned to each representation of the transition distribution. Furthermore, we use a full-rank covariance matrix, instead of a less expressive diagonal covariance.
Related work also includes Hidden Parameter MDPs (HiP-MDPs) which assume the agent knows what it does not know. Specifically, it does not know the value of some critical environment variables and focuses on estimating these [doshi2016hidden]. [zhang2020learning] learn state abstractions to enable multi-task RL. However, they assume the latent parameter is fixed for the duration of the episode, while in our case we are interested in detecting and accommodating to changes in the latent parameters during the episode. [killian2017robust] extend this framework for robust and efficient transfer learning to subtle variations among tasks using Bayesian Neural Networks. [costen2023planning] develop a method to express the transition probabilities in closed-form as a polynomial of the latent parameters, as well as maintaining a closed-form belief distribution of the latent parameters. They present promising results in two simulation tasks. In our case, the transition model’s input is augmented with the symbol of the detected latent representation, since we do not assume prior-knowledge about the variables that the agent will require to obtain a better model of the dynamics at deployment time.